the user is charge p 300 monthly for a particular mobile plan, which includes 100 free text messages.
messages in excess of 100 are charged p 1 each.
represent the amount t a consumer pays each month as a function of the number of messages m sent in a month

Answer by Theo(13342)

this would be a partial function or what might be called a split function.

t = 300 for x <= 100

t = 300 + 1 * (x-100) for x > 100.

as long as the number of messages is less than or equal to 100, the first formula of t = 300 is used.

if the number of messages exceeds 100, then the second formula of t = 300 + 1 * (x - 100) is used.

assume the customer used 150 messages.
since the number of messages is greater than 100, the second formula kicks in.
the monthly charge is 300 + 1 * (150 - 100) = 300 + 1 * 50 = 300 + 50 = 350.

assume the customer used 90 messages.
since the number of messages is less than or equal to 100, the first formula kicks in.
the monthly charge is 300.